OpenStudy (anonymous):

The equation s=p(1+r)^t can be used to model a person’s future income. S = future salary, p = current salary, r = rate of increase, and t = time in years. Susan’s salary increases by 3% each year. If she starts at $43,000 per year, what will her salary be in 5 years?

OpenStudy (anonymous):

@Paul2R

OpenStudy (anonymous):

thanks..so this is basically a plug and chug

OpenStudy (anonymous):

what is r in the problem?

OpenStudy (anonymous):

The r is 1 + 0.03

OpenStudy (anonymous):

so what are you saying?

OpenStudy (anonymous):

yup..that is becuase the salary is increasing..so it would be current salary plus the extra 3%

OpenStudy (anonymous):

@coolsday & @Paul2R

OpenStudy (anonymous):

now what is P?

OpenStudy (anonymous):

idk @Paul2R

OpenStudy (anonymous):

what salary are you starting at will be P. r is the rate of increase, and t is the amount in years.

OpenStudy (anonymous):

wht is the girls current salary?

OpenStudy (anonymous):

ohhh okay hold on lol @Paul2R

OpenStudy (anonymous):

$46,750? @Paul2R

OpenStudy (anonymous):

is that right @coolsday

OpenStudy (anonymous):

P = 43,000 r = 1.03 t = 5 You want to find s s = P(1+r)^t s=43000(1.03)^5 can you solve for x?

OpenStudy (anonymous):

you are close, but that is not right

OpenStudy (anonymous):

49,850? @coolsday

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank youuuu! @coolsday

OpenStudy (anonymous):

np